Treasury Yields. What is the yield to maturity for Treasury strips with a quoted price of 87.45 with 4.5 years to maturity? Assume $100,000 par. What is the price of a $1 million Treasury Bill with 156 days to maturity and a discount yield of 2.15%?
Answer #1

1: Using financial calculator
Input: FV= 100000

PV = -87.45%*100000 = 87450

N = 4.5

Solve for I/Y as 3.02

Hence YTM = 3.02%

2: Discount yield = (F-P)/F * (360/T)

0.0215 = (100000-P)*(360/156) /100000

2150= (100000-P)*2.30769

931.67 = (100000-P)

P = 99068.33